### v1

@@ 3 Measurements to reveal quantumness
In this section, we address the first two previous points. We begin by emphasizing that the Bell inequality provides answers to very specific questions (as will be addressed in the next section) and requires extreme caution when applied to verify the quantumness of correlations [19].
@@
In the case of classical entanglement, the vector space is partitioned in a fundamentally different manner (as discussed in subsection 2.2). There is only one excitation present, which can be assigned to either of the partitions spanned by available mode functions. Consider the same prototypic example as in Eq. (4):
@@ 4 Locality and non-locality of non-separable states
As strongly emphasized in [19], “the main deviation of classical light models from quantum theory is not only in the states but also in the description of measurement procedures." Khrennikov argues strongly against using quantum non-locality to differentiate between quantum and classical correlations, attributing the crucial role to measurement procedures [19], which aligns well with our arguments. Importantly, the conventional view is that quantum theory is local, but the assumption of realism is incorrect. It is quite straightforward to argue for this. The principle of locality dictates that interactions propagate with finite velocity. However, when the measurement outcome on $A$ determines the measurement outcome on $B$ in a non-separable state, it is not about interaction between these two subsystems. A system in a non-separable state is in a superposition of two possibilities (in the bipartite case): for the state of Eq. (6), the system is simultaneously in state $|0\rangle_{A}|1\rangle_{B}$ and $|1\rangle_{A}|0\rangle_{B}$; it is not in a definite, ‘pre-existing’ state. The distinct state emerges only upon measurement and is ’decided’ at random in the process of measurement with a given probability. This is a clear demonstration of the violation of the principle of reality, which states that properties of objects are real and exist independent of our observations of them.
Importantly, the conventional view is that quantum theory is local, yet the assumption of realism is incorrect. This argument is quite straightforward. The principle of locality stipulates that interactions propagate with finite velocity. However, when the measurement outcome on $A$ determines the outcome on $B$ in a non-separable states, it actually happens not due to the interaction between these two subsystems. Instead, a system in a non-separable state is in a superposition of two (in bi-partite systems) possibilities: For the state of Eq. (6) the system is simultaneously in the state $|0\rangle_{A}|1\rangle_{B}$ and $|1\rangle_{A}|0\rangle_{B}$, it is not in a definite, ‘pre-existing’ state. The distinct state emerges only upon measurement and is ‘decided’ at random in the process of measurement with a given probability.
@@
Figure 4: Principle experimental schemes for construction of Bell inequality for verification of polarization correlations $f(\theta,\theta^{\prime})$. Paramerization by $\theta$ happens either as part of the measurement process, or as part of the state preparation process, as indicated by dashed boxes. (Top Left) In case of LHV assumption, the polarization is predetermined at the source and two measurements are performed parametrized by $\theta$; (Bottom Left) In case of quantum entanglement, the polarization is not defined at the source and is determined in the measurement process, at the detector, when two measurements are performed parametrized by $\theta$; (Right) In case of classical non-separability, light at the source is prepared in a certain polarization state parameterized by $\theta$. A single, parametrized by $\theta^{\prime}$ measurement is performed.
@@
Let us depart for the moment from any specific experimental details on how it has been implemented, but let us focus on the principal idea of an experimental setting used to construct the inequality [13, 12]. The statistical analysis required begins with performing two independent measurements on two parts of the Hilbert space, these measurements being parameterized by some parameter $\theta$. For example, in the case of polarization entanglement of Eq. (1), the measurement process will include a polarizer, set in front of a detector and rotated at a certain angle $\theta$. In other words, the measurement will include a detector and a ‘filter’ set for $\theta$, as discussed previously. The inequality is constructed by combining probabilities of measurement outcomes obtained at detector 1 plus filter $\theta$ and detector 2 plus filter $\theta^{\prime}$ for particular values of $\theta$ and $\theta^{\prime}$ (for details, we refer the interested reader to [13, 12]).
@@ 5 Conclusions
With great pleasure, we dedicate this article to Rodney Loudon. His legacy continues to inspire scientists and pave the way for further exploration in the realm of quantum optics. Over the years, we have benefited from questions, suggestions, criticism, and advice from many colleagues. Particular gratitude for help in various ways goes to Andrea Aiello, Gunnar Björk, Bob Boyd, Joe Eberly, Robert Fickler, Elisabeth Giacobino, Aaron Goldberg, Ebrahim Karimi, Antonio Z. Khoury, Andrei Khrennikov, and Andrei Klimov.

### v3

@@ 3 Measurements to reveal quantumness
In this section, we address the first two previous points. We begin by emphasizing that the Bell inequality provides answers to very specific questions (as will be addressed in the next section) and requires extreme caution when applied to verify the quantumness of correlations [22]. Therefore we take here a measurement-based, operational approach which has common points with earlier work of [16, 17] and recent contributions by [22, 23]. The approach of [16, 17] is based on observable induced choice of Hilbert space partitioning as already discussed in section 2. The most recent work of Khrennikov and Basieva [23] brings the idea of observable induced partitioning [17] even further and aims to decouple the notion of entanglement from the tensor product structure and associated non-separability. Their work emphasizes the role of measurement process even more: Khrennikov and Basieva develop the probabilistic formalization of the Schrödinger’s concept of entanglement without reference to any complex vector spaces. Entanglement is defined as dependence of observables $A$ and $B$; relevant constraints are derived based on the quantum conditional probability calculus. In our approach and the approaches referred above, choice of observables, operational access to degrees of freedom, and hence a measurement process as such, step forward as pivotal tools to assess quantum entanglement presence.
@@
In the case of classical entanglement, the vector space is partitioned in a fundamentally different manner (as discussed in subsection 2.2). There is only one excitation present, which can be assigned to either of the partitions spanned by available mode functions. Consider the same prototypical example as in Eq. (8):
@@ 4 Locality and non-locality of non-separable states
In early 2000s, the discussion of the former aspect (is entanglement primarily about non-locality) has already led to a counter-intuitive conclusion that entanglement and non-locality are different resources. As an example, a thorough investigation of incommensurability between entanglement and non-locality is presented in [24]. We return to this question a bit later in this section in the context of the measurement-based approach taken.
@@
As strongly emphasized in [22], “the main deviation of classical light models from quantum theory is not only in the states but also in the description of measurement procedures." Khrennikov argues strongly against using quantum non-locality to differentiate between quantum and classical correlations, attributing the crucial role to measurement procedures [22], which aligns well with our arguments. This “measurement-based” approach provides a key to interpreting the violation of the local realism. Importantly, the conventional view is that quantum theory is local, yet the assumption of realism is incorrect. This argument is quite straightforward. The principle of locality stipulates that interactions propagate with finite velocity. However, when the measurement outcome on $A$ determines the outcome on $B$ in a non-separable states, it actually happens not due to the interaction between these two subsystems. Instead, a system in a non-separable state is in a superposition of two (in bi-partite systems) possibilities: For the state of Eq. (10) the system is simultaneously in the state $|0\rangle_{A}|1\rangle_{B}$ and $|1\rangle_{A}|0\rangle_{B}$, it is not in a definite, ‘pre-existing’ state. The distinct state emerges only upon measurement and is ‘decided’ at random in the process of measurement with a given probability.
@@
Figure 4: Principle experimental schemes for construction of Bell inequality for verification of polarization correlations $f(\theta,\theta^{\prime})$. Parameterization by $\theta$ happens either as part of the measurement process, or as part of the state preparation process, as indicated by dashed boxes. (Top Left) In case of LHV assumption, the polarization is predetermined at the source and two measurements are performed parameterized by $\theta$; (Bottom Left) In case of quantum entanglement, the polarization is not defined at the source and is determined in the measurement process, at the detector, when two measurements are performed parameterized by $\theta$; (Right) In case of classical non-separability, light at the source is prepared in a certain polarization state parameterized by $\theta$. A single, parameterized by $\theta^{\prime}$ measurement is performed.
@@
Let us depart for the moment from any specific experimental details on how it has been implemented, but let us focus on the principal idea of an experimental setting used to construct the inequality [13, 12]. The statistical analysis required begins with performing two independent measurements on two parts of the Hilbert space, these measurements being parameterized by some parameter $\theta$. For example, in the case of polarization entanglement of Eq. (2), the measurement process will include a polariser, set in front of a detector and rotated at a certain angle $\theta$. In other words, the measurement will include a detector and a ‘filter’ set for $\theta$, as discussed previously. The inequality is constructed by combining probabilities of measurement outcomes obtained at detector 1 plus filter $\theta$ and detector 2 plus filter $\theta^{\prime}$ for particular values of $\theta$ and $\theta^{\prime}$ (for details, we refer the interested reader to [13, 12]).
@@ 5 Conclusions
In the same vein, this increased parameter space has been used in the pervasive field of indefinite causal order, where one can think of using one degree of freedom as control for an operation that is acting on another degree of freedom. Incorporating this control in particular tasks leads to advantages in computation [33], communication [34], and metrology [35, 36], many of which have been experimentally realized [37, 38, 39, 40, 41].
We propose an operational distinction to clearly identify quantum entanglement and emphasize that two measurement are required for this identification. The projective quantum measurement processes by themselves remains a mystery: trying to find, where in the chain of events the projection happens is a moving target, and the „speed“ at which the projection happens seems to be solely determined by the response time of the detector.
@@
With great pleasure, we dedicate this article to Rodney Loudon. His legacy continues to inspire scientists and to pave the way for further exploration in the realm of quantum optics. Over the years, we have benefited from questions, suggestions, criticism, and advice from many colleagues. Particular gratitude for help in various ways goes to Andrea Aiello, Gunnar Björk, Bob Boyd, Joe Eberly, Robert Fickler, Elisabeth Giacobino, Aaron Goldberg, Ebrahim Karimi, Antonio Z. Khoury, Andrei Khrennikov, and Andrei Klimov.